OpenStudy (anonymous):

two objects of same mass in a head on collision both 15kg, Object 1 Vi - 1.5 m/sec and Vf- 2m/sec what is object 2 Vf and Vi

OpenStudy (anonymous):

collision is elastic

OpenStudy (anonymous):

so you'll want to use the conservation of momentum equation \[m_1*v_1 + m_2*v_2 = m_1*v_1 + m_2*v_2\] where the left of the = sign is 'initial' and the right of the equal sign is 'final' so with the numbers we have, our equation would look like: \[15_1(-1.5)_1 + (15_2)v_2 = (15_1)2_1 + 15_2v_2\] so we are missing Object 2 velocity initial and velocity final. now since the masses are the same for both object 1 and 2, they will have the same velocity after the collision! so object 2 final velocity is the same as object 1 final velocity so, 2 m/s = FinalV_1 and FinalV_1 = FinalV_2 can you see how to finish the rest?